# Unifying Specialized Visual Encoders for Video Language Models

### v1

@@ 6 Conclusion
Previous VideoLLMs have been limited to relying on a single visual model for feature extraction, which leads to limited understanding capabilities of vastly different video tasks. In our work, we break this paradigm and explore various fusion strategies for combining information from multiple visual experts to generate a representation that can leverage the capabilities of different video encoders. We find that our multi-encoder feature fusion is able to outperform comparable methods by up to 3.79% on video reasoning benchmarks. We show that the method can obtain better performance than the best-performing single-encoder model with minimal computational overhead. Finally, we quantitatively and qualitatively observe the skill specializations our model learns on an MCQ format of Something-Something v2, which confirms both that encoders can be specialized and that our model captures both axes of knowledge. Our paper proposes some initial steps in rethinking how we approach the use of multiple encoders. We hope that this inspires others to also consider this problem as potentially another direction for scaling and improving their VideoLLMs.
Acknowledgements: This work is supported by the National Science Foundation under Grant No. 2107048 and the Princeton First Year Fellowship to TZ. We also thank Allison Chen and William Yang for detailed comments and feedback.
@@ Appendix A Appendix / supplemental material
A.1 Limitations and Broader Impact.
Our works are based on LLaMA-2 7B model (Touvron et al., 2023b), and as with many other VideoLLM models, the performance of our method is hugely dependent on the capabilities of the LLM model, and better-performing models often demand significantly more computation. MERV requires running multiple encoders, which can be computationally intensive and can lead to out-of-memory errors in limited resource settings. While FSDP (Zhao et al., 2023) allows us to easily and effectively train larger models across multiple GPUs than would otherwise be possible, its generality also makes it difficult for us to design tailored sharding strategies that would maximize the performance of our model. However, with future improvements to data parallelism, our model can still benefit greatly and run even more efficiently. Also, despite the improved speeds, there is still an upper bound for what constitutes a reasonable training time that still allows us to test many of our design assumptions, which limits the scale and number of experiments we can run.
While we show that our method can successfully leverage information from different visual encoders, nevertheless if the encoders themselves are limited in video understanding capability, MERV cannot fully compensate for that. Our work aims to facilitate video understanding, which can lead to positive social impacts such as a video captioning model for low-vision users, automatic detection of medical emergencies, or better self-driving cars. It can also lead to negative social impacts like easy surveillance by the authorities, and human-like internet bots being used for scamming purposes. We follow the same safeguards implemented by the original authors of the datasets, the visual models, and the LLM models. We have not put any additional safeguards ourselves.
A.2 Qualitative Results
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 7: Samples of MERV in video understanding
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 8: MERV examples. MERV tend to show improved understanding in temporal-heavy videos as in Something-Something v2 dataset (Goyal et al., 2017) (Top Row), while retaining the performance on scenic understanding, seen from popular video benchmarks (Xu et al., 2017; Yu et al., 2019) (Bottom Row).
Appendix B Training Detail
B.0.1 Baseline Encoder and LLM Details
@@
LLaMA-Adapter (Zhang et al., 2023b) CLIP (Radford et al., 2021) LLaMA-1 7B (Touvron et al., 2023a)
Video-LLaMA (Zhang et al., 2023a) ViT-G (EVA-CLIP) (Sun et al., 2023b) + BLIP-2 Q-Former (Li et al., 2023b) Vicuna-7B v0 (Chiang et al., 2023)
Video-ChatGPT (Maaz et al., 2023) CLIP (Radford et al., 2021) Vicuna-7B v1.1 (Chiang et al., 2023)
SeViLA (Yu et al., 2024) ViT-G (EVA-CLIP) (Sun et al., 2023b) + BLIP-2 Q-Former (Li et al., 2023b) FlanT5-XL (3B) (Chung et al., 2024)
LLaMA-VID-7B (Li et al., 2023e) EVA-G (Fang et al., 2023) Vicuna-7B v1.5 (Chiang et al., 2023)
LLaMA-VID-13B (Li et al., 2023e) EVA-G (Fang et al., 2023) Vicuna-13B v1.5 (Chiang et al., 2023)
Video-LLaVA* (Lin et al., 2023a) LanguageBind (Zhu et al., 2023) Vicuna-7B v1.5 (Chiang et al., 2023)
Table 3: Visual Encoder and LLM Information.
B.0.2 MERV Encoder Details
@@
LanguageBind (Zhu et al., 2023) ViT-L/14 Video+Language VIDAL-10M, five-modal video examples Contrastive
@@
Here, we detail the visual encoder details, LLM, and the training objectives. We plan to release the code for the camera-ready version of the paper.
@@
DINOv2 As DINOv2 is an image-model, we get embedding per frame, and concatenate them to be a video embedding. We use $\texttt{ViT}_{\texttt{Large}}$ model, pre-trained on LVD-142M dataset, and take the penultimate layer for the embeddings. Specifically, we use timm’s vit_large_patch14_reg4_dinov2l̇vd142m
@@
We also considered multiple other options for encoders, such as CLIP-ViP (Xue et al., 2022) for our video-language contrastive expert, V-JEPA (Bardes et al., 2024) and Hiera (Ryali et al., 2023) for our pure video model, and CLIP (Radford et al., 2021) for our image-language contrastive expert, but found that our choices performed better overall.
B.1 Detailed Experimental Results.
@@
Methods MSVD-QA MSRVTT-QA TGIF-QA Perception ActivityNet-QA
@@
We also provide the full numbers for our ablations in Sections 4.2.1 and 4.2.2.
@@
Table 7: Effect of Each Encoder. Detailed results of Figure 4a
@@
Table 9: Performance on WH-words. Detailed experimental results of Figure 5
B.2 Something Something v2 Details
@@
B.2.1 Something-Something V2 - OpenEnded.
Additionally, we evaluate Something-Something V2 as an open ended QA task, where the question is "How is the object in the video being interacted with?", and the answer is expected to be similar to the class label. We use Video-ChatGPT (Maaz et al., 2023)’s LLM evaluation for validating the VideoLLMs’ output. Table 10 tabulates the results.
B.2.2 Something-Something V2 - Temporal.
@@
Appendix C Additional Analysis
C.1 Attention Weights

### v2

@@ 6 Conclusion
Previous VideoLLMs have been limited to relying on a single visual model for feature extraction, which leads to limited understanding capabilities of vastly different video tasks. In our work, we break this paradigm and explore various fusion strategies for combining information from multiple visual experts to generate a representation that can leverage the capabilities of different video encoders. We find that our multi-encoder feature fusion is able to outperform comparable methods by up to 4.62% on video reasoning benchmarks. We show that the method can obtain better performance than the best-performing single-encoder model with minimal computational overhead. Finally, we quantitatively and qualitatively observe the skill specializations our model learns on an MCQ format of Something-Something v2, which confirms both that encoders can be specialized and that our model captures both axes of knowledge. Our paper proposes some initial steps in rethinking how we approach the use of multiple encoders. We hope that this inspires others to also consider this problem as another direction for scaling and improving their VideoLLMs.
Acknowledgements
This work is supported by the National Science Foundation under Grant No. 2107048 and the Princeton First Year Fellowship to TZ. We also thank Allison Chen and William Yang for detailed comments and feedback.
Impact Statement
Our work aims to facilitate video understanding, which can lead to positive social impacts such as a video captioning model for low-vision users, automatic detection of medical emergencies, or better self-driving cars. It can also lead to negative social impacts like easy surveillance by the authorities, and human-like internet bots being used for scamming purposes. We follow the same safeguards implemented by the original authors of the datasets, the visual models, and the LLM models. We have not put any additional safeguards ourselves.
@@ Appendix A Appendix / Supplemental Material
A.1 Limitations
Our work is based on the LLM, LLaMA-2 7B model (Touvron et al., 2023b), and as with many other VideoLLM models, the performance of our method is hugely dependent on the capabilities of the LLM model, and better-performing models often demand significantly more computation. MERV requires an LLM and running multiple encoders, which can be computationally intensive and can lead to out-of-memory errors in resource-limited settings (while our efficient implementation alleviates the issue). While FSDP (Zhao et al., 2023) allows us to easily and effectively train larger models across multiple GPUs than would otherwise be possible, its generality also makes it difficult for us to design tailored sharding strategies that would maximize the performance of our model. However, with future improvements to data parallelism, our model can still benefit greatly and run even more efficiently. Also, despite the improved speeds, there is still an upper bound for what constitutes a reasonable training time that still allows us to test many of our design assumptions, which limits the scale and number of experiments we can run. While we show that our method can successfully leverage information from different visual encoders, nevertheless if the encoders themselves are limited in video understanding capability, MERV cannot fully compensate for that.
A.2 Model Details
A.2.1 Baseline Encoder and LLM Details
Table 3: Visual Encoder and LLM Information.
@@ A.2.1 Baseline Encoder and LLM Details
LLaMA-Adapter (Zhang et al., 2024b) CLIP (Radford et al., 2021) LLaMA-1 7B (Touvron et al., 2023a)
Video-LLaMA (Zhang et al., 2023) ViT-G (EVA-CLIP) (Sun et al., 2023b) + BLIP-2 Q-Former (Li et al., 2023b) Vicuna-7B v0 (Chiang et al., 2023)
Video-ChatGPT (Maaz et al., 2024) CLIP (Radford et al., 2021) Vicuna-7B v1.1 (Chiang et al., 2023)
SeViLA (Yu et al., 2024b) ViT-G (EVA-CLIP) (Sun et al., 2023b) + BLIP-2 Q-Former (Li et al., 2023b) FlanT5-XL (3B) (Chung et al., 2024)
LLaMA-VID-7B (Li et al., 2024b) EVA-G (Fang et al., 2023) Vicuna-7B v1.5 (Chiang et al., 2023)
LLaMA-VID-13B (Li et al., 2024b) EVA-G (Fang et al., 2023) Vicuna-13B v1.5 (Chiang et al., 2023)
Video-LLaVA* (Lin et al., 2024) LanguageBind (Zhu et al., 2024a) Vicuna-7B v1.5 (Chiang et al., 2023)
A.2.2 MERV Encoder Details
@@
LanguageBind (Zhu et al., 2024a) ViT-L/14 Video+Language VIDAL-10M, five-modal video examples Contrastive
@@
Here, we detail the visual encoder details, LLM, and the training objectives. We plan to release our full code including training and evaluation as well as all model checkpoints for the camera-ready version of the paper.
@@
DINOv2 As DINOv2 is an image-model, we get embedding per frame, and concatenate them to be a video embedding. We use $\texttt{ViT}_{\texttt{Large}}$ model, pre-trained on LVD-142M dataset, and take the penultimate layer for the embeddings. Specifically, we use timm’s vit_large_patch14_reg4_dinov2.lvd142m
@@
We also considered multiple other options for encoders, such as CLIP-ViP (Xue et al., 2023) for our video-language contrastive expert, V-JEPA (Bardes et al., 2024) and Hiera (Ryali et al., 2023) for our pure video model, and CLIP (Radford et al., 2021) for our image-language contrastive expert, but found that our choices performed better overall.
A.3 Detailed Experimental Results
@@
Methods MSVD-QA MSRVTT-QA TGIF-QA Perception Test ActivityNet-QA
@@
We also provide the full metric numbers in Table 6a6b and 6c for our ablations that are described in Sections 4.2.1 and 4.2.2.
@@
Table 7: Effect of Each Encoder. Detailed results of Figure 4a.
@@
Table 9: Performance on WH-words. Detailed experimental results of Figure 6.
@@
Note on TGIF-QA
Video-ChatGPT’s and Video-LLaVA’s author-reported numbers on TGIF are incomparable as they were on a subset of the dataset. See https://github.com/PKU-YuanGroup/Video-LLaVA/issues/37.
A.4 Something-Something v2 Details
@@ A.4.1 Something-Something v2 - OpenEnded
Additionally, we evaluate Something-Something V2 as an open ended QA task, where the question is ”How is the object in the video being interacted with?”, and the answer is expected to be similar to the class label. We use Video-ChatGPT (Maaz et al., 2024)’s LLM evaluation for validating the VideoLLMs’ output. Table 10 tabulates the results. We see similar conclusions as with MCQ where ViViT excels at temporal subset than other single-encoder LLMs, but failing to match the performance in full dataset. Nonetheless, our method shows the best performance overall.
A.4.2 Something-Something v2 - Temporal
@@
A.5 Additional Experiments and Analyses
A.5.1 Adding More Encoders
Additionally, we test with one more encoder Hiera (Ryali et al., 2023). The checkpoint we use was pretrained using the Masked Autoencoder (MAE) self-supervised learning technique on Kinetics 400 (K400), and then finetuned on K400 using supervised learning.
Table 11: Performance with Hiera (Ryali et al., 2023).
Methods MSVD-QA MSRVTT-QA TGIF-QA Perception ActivityNet-QA
Acc Score Acc Score Acc Score Acc Acc Score
VideoLLaVA 67.74 3.69 56.90 3.18 47.99 3.17 47.08 3.27 44.22
MERV 70.97 3.76 59.03 3.25 51.1 3.26 50.87 3.34 46.21
MERV, ViViT replaced with Hiera 69.68 3.74 57.64 3.22 50.38 3.24 50.24 3.34 47.50
MERV + Hiera 69.67 3.72 58.26 3.23 50.32 3.22 51.23 3.34 46.23
ViViT Single Encoder LLM 59.95 3.43 51.81 3.05 38.10 2.84 43.98 3.16 40.20
Hiera Single Encoder LLM 55.38 3.28 49.21 2.95 36.02 2.76 44.01 3.15 40.20
The results are presented in the Table 11, with the best result in bold. Replacing ViViT with Hiera demonstrates improvements on the fine-grained spatio-temporal reasoning benchmark, Perception Test, with accuracy gain of 1.29%. Similarly, adding Hiera yields an improvement on ActivityNet, achieving a 0.36% increase in accuracy. However, on other benchmarks, the original MERV remains the strongest model. Overall, we observe no significant performance improvement when training with Hiera, which aligns with expectations, since Hiera is under the same paradigm as ViViT, functioning as a temporal expert trained on short videos. We also hypothesize that Hiera is more sensitive to the temporal stride than ViViT, as ViViT can reasonably deduce motion from uniformly sampled frames. We expect performance to improve if we incorporate encoders trained on different paradigms and data sources or process a much greater number of frames simultaneously, which we will leave for future work.
A.5.2 Attention Weights
@@
A.6 More Qualitative Results
We present additional qualitative results in Figure 12 and Figure 13.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 12: Samples of MERV in video understanding.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 13: More qualitative results. MERV tends to show improved understanding in temporal-heavy videos as in Something-Something v2 dataset (Goyal et al., 2017) (Top Row), while retaining the performance on scenic understanding, seen from popular video benchmarks (Xu et al., 2017; Yu et al., 2019) (Bottom Row).